OpenStudy (anonymous):

Jane is painting her room violet. The 15 gallon bucket of purple paint that she currently has is 30% red paint and 70% blue paint. This shade of violet is more red, and so Molly will need paint that is 40% red paint, and thus 60% blue paint. How many gallons of red paint should she add to her current bucket in order to be able to paint her room violet?

OpenStudy (anonymous):

@DangerousJesse Are you still there? lol jw :)

OpenStudy (dangerousjesse):

\(\frac{3}{7}=\frac{9}{21}\) \(\frac{4}{6}=\frac{6}{9}\) So that's 20% more red paint needed \(15\times 1.2= 3\) 3 gallons

OpenStudy (dangerousjesse):

Sorry, mathjax froze :P

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

I actually understand it now and will redo the problem on my own but this helped me a lot! @DangerousJesse